A 100m^2 factory flat roof feeds all the water collected to a rainwater tank. If there is 1mm of rainfall, how many litres of water go into the tank?
Answer by math_tutor2020(3817)


mL = milliliters
1 mL = 1 cubic cm
Think of a 1 cm by 1 cm by 1 cm cube.

1 meter = 100 cm
(1 meter)^2 = (100 cm)^2
1 square meter = 10,000 square cm
100*(1 square meter) = 100*(10,000 square cm)
100 square meters = 1,000,000 square cm

10 mm = 1 cm
1 mm = 1/10 of a cm
1 mm = 0.1 cm

Volume of rectangular prism. Think of a rectangular slab.
volume = (area of the base)*(height)
volume = (1,000,000 square cm)*(1 mm)
volume = (1,000,000 square cm)*(0.1 cm)
volume = 100,000 cubic cm
volume = 100,000 mL

1000 mL = 1 liter
100*(1000 mL) = 100*(1 liter)
100,000 mL = 100 liters


Answer: 100 liters